I'm stumped. I tried multiplying by 3 to isolate the 2x, and tried combining fractions then dividing, but none of my attempts have been right. Can anyone help please?
Solve the inequality. Write the solution set in interval notation.
-2/3 >= 2x/3 - 3/4


Found 2 solutions by stanbon, coyote:
Answer by stanbon(75887)

-2/3 >= 2x/3 - 3/4
Multiply thru by 12 to get rid of the denominators:
-8 >= 8x -9
1 <= 8x
x >= 1/8
-----------
Interval notation: [1/8,inf)
============
Cheers,
Stan H.

Answer by coyote(32)

-2/3 >= 2x/3 - 3/4
Multiply thru by 12 to get rid of the denominators:
-8 >= 8x -9
1 <= 8x
x >= 1/8
-----------
Interval notation: [1/8,inf)

Almost, but you helped immensely when I realized the one >= was just reversed!
Thank you!
-2/3 >= 2x/3 - 3/4
Multiply thru by 12 to get rid of the denominators:
-8 >= 8x -9
1 >= 8x
x <= 1/8
-----------
Interval notation: [inf,1/8)